Find the roots of 5x^6+4x^5+4x^3-61x^2-48x+12=0
1
Expert's answer
2012-01-05T11:30:35-0500
5x^6+4x^5+4x^4+4x^3-61x^2-48x+12=0
we check all numbers that divide 12/5 so
we get -1 and 1/5 roots of equation
then we divide polinomial on (x+1)(5x-1)
we get x^4+x^2-12=0 we see that it is biquadratic equation so x^2=-4, x^2=3 and
get x=+ or - sqrt(3), x=+ or - 2i